i need help solving this ASAP. Please help me convert this to the slope-intercept form!
y+10=-4/5(x-5)

Answer by jim_thompson5910(35256)

y+10=-4/5(x-5)

y+10=-4/5x+(-4/5)(-5)

y+10=-4/5x+4

y=-4/5x+4-10

y=-4/5x-6


If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim